# Similarity Memory Prior is All You Need for Medical Image Segmentation

### v1

@@ Abstract
In recent years, it has been found that “grandmother cells” in the primary visual cortex (V1) of macaques can directly recognize visual input with complex shapes. This inspires us to examine the value of these cells in promoting the research of medical image segmentation. In this paper, we design a Similarity Memory Prior Network (Sim-MPNet) for medical image segmentation. Specifically, we propose a Dynamic Memory Weights–Loss Attention (DMW-LA), which matches and remembers the category features of specific lesions or organs in medical images through the similarity memory prior in the prototype memory bank, thus helping the network to learn subtle texture changes between categories. DMW-LA also dynamically updates the similarity memory prior in reverse through Weight-Loss Dynamic (W-LD) update strategy, effectively assisting the network directly extract category features. In addition, we propose the Double-Similarity Global Internal Enhancement Module (DS-GIM) to deeply explore the internal differences in the feature distribution of input data through cosine similarity and euclidean distance. Extensive experiments on four public datasets show that Sim-MPNet has better segmentation performance than other state-of-the-art methods. Our code is available on https://github.com/vpsg-research/Sim-MPNet.
@@ 3.2 Dynamic Memory Weight-Loss Attention
In the above formula, $\alpha$ and $\beta$ are decay factors, which are set to 0.5, and $\theta$ = $M$. In addition, we set $\frac{M}{4}\leq K\leq\frac{3M}{4}$ to avoid $K$ falling into extreme situations and ensure stable training. In this way, we can not only minimize the loss of information but also ensure the semantic clarity and representation ability of the similarity memory prior.
@@ 3.3 Double-Similarity Global Internal Enhancement Module
Table 2: Performance on Synapse. P and F represent Parameters and FLOPs respectively. ${T}_{i}$ is the average inference efficiency of a single image in milliseconds. DSC are reported for individual organs. GB, KL, KR, PC, SP and SM represent gallbladder, left kidney, right kidney, pancreas, spleen and stomach respectively.
@@ 4.1 Datasets
Figure 7: Visualization of semantic clusters (similarity memory priors) using DMW-LA on Synapse, ACDC, and SegPC-2021. We calculate the distribution of similarity memory prior in the prototype memory bank, and highlight their receptive fields with white boxes. We further group the similarity memory priors by clusters (columns). For example, for “Cluster 2”, all white boxes are receptive fields formed by Cluster 2.
@@ 4.2 Implementation Details and Evaluation Metrics
In all experiments, the hyperparameters of DS-GIM (decay coefficient) and DMW-LA (linear function) remained unchanged. For different datasets, just set k in K-means as the number of categories in the dataset. The input resolutions of two encoders are set to 224$\times$224 and 256$\times$256 respectively. In addition, we train the model using an AdamW optimizer with a learning rate of 0.0001, and a weight decay of 1e-4. The model is trained on a single RTX 3090 GPU with a batch size 20 using Dice ($L_{dice}$) and Cross-entropy ($L_{ce}$) losses for 300 epochs. The specific proportion of these two losses is $Loss=0.7\cdot L_{dice}+0.3\cdot L_{ce}$. Finally, all experiments are performed in the Python 3.10.0 and Torch 2.3.0 environment. Only use random rotation and flipping as data augmentation. We choose the Dice score (DSC) and 95% Hausdorff distance (HD95) as evaluation metrics. The HD95 represents the ${95}^{th}$ percentile distance between the boundaries of the model prediction graph and real labels. See Appendix 5 for more details.
@@ 4.3 Comparison with SOTA Methods
textbfVisual Comparison and Analysis. Figure 6 shows the visualizations of Sim-MPNet and other networks on Synapse and SegPC-2021. Firstly, the segmentation results of Sim-MPNet are very close to the ground truth in Synapse. Especially when dealing with organs like the Pancreas, which have complex contours and rich details, Sim-MPNet shows better segmentation performance compared to other networks. Secondly, Sim-MPNet still obtains the best segmentation results for images containing multiple cells in SegPC-2021. Although false positives and false negatives may also occur, the incidence of these errors is significantly reduced compared to other networks. In the process of medical image segmentation, we assign a semantic cluster (similarity memory prior) to the segmentation target in each image to extract key categorical features. Figure 7 shows the semantic cluster receptive field of DMW-LA on three datasets. We can observe that different clusters encode different semantic concepts. In the Synapse dataset, Cluster 2, Cluster 4 and Cluster 6 tend to focus on the region where the organs are located. In the ACDC dataset, Cluster 1 tends to focus on the left ventricle, while Cluster 2 focuses more on the right ventricle. In the SegPC-2021 dataset, Cluster 1 tends to focus on cytoplasm, while Cluster 3 focuses more on tissue fluid and other cells in the background. This distribution indicates that when processing images, each semantic cluster focuses on specific category features, thereby helping the model identify key structures in medical images.
@@
Table 3: Ablation study of DMW-LA on four datasets. We use different encoders to verify their effectiveness separately. “—” means not to use.

### v2

@@ Abstract
In recent years, it has been found that “grandmother cells” in the primary visual cortex (V1) of macaques can directly recognize visual input with complex shapes. This inspires us to examine the value of these cells in promoting the research of medical image segmentation. In this paper, we design a Similarity Memory Prior Network (Sim-MPNet) for medical image segmentation. Specifically, we propose a Dynamic Memory Weights-Loss Attention (DMW-LA), which matches and remembers the category features of specific lesions or organs in medical images through the similarity memory prior in the prototype memory bank, thus helping the network to learn subtle texture changes between categories. DMW-LA also dynamically updates the similarity memory prior in reverse through Weight-Loss Dynamic (W-LD) update strategy, effectively assisting the network directly extract category features. In addition, we propose the Double-Similarity Global Internal Enhancement Module (DS-GIM) to deeply explore the internal differences in the feature distribution of input data through cosine similarity and euclidean distance. Extensive experiments on four public datasets show that Sim-MPNet has better segmentation performance than other state-of-the-art methods. Our code is available on https://github.com/vpsg-research/Sim-MPNet.
@@ 3.2 Dynamic Memory Weight-Loss Attention
In the above formula, $\alpha$ and $\beta$ are decay factors, which are set to 0.5, and $\theta$ = $M$. In addition, we set $\frac{M}{4}\leq K\leq\frac{3M}{4}$ to avoid $K$ falling into extreme situations and ensure stable training, see appendix A. In this way, we can not only minimize the loss of information but also ensure the semantic clarity and representation ability of the similarity memory prior.
@@ 3.3 Double-Similarity Global Internal Enhancement Module
Table 2: Performance on Synapse. P and F represent Parameters and FLOPs respectively. FLOPs of all the methods are reported for 224$\times$224 inputs, except Sim-MPNet (dual encoder: 224$\times$224, 256$\times$256). ${T}_{i}$ is the average inference efficiency of a single image in milliseconds. DSC are reported for individual organs. GB, KL, KR, PC, SP and SM represent gallbladder, left kidney, right kidney, pancreas, spleen and stomach respectively.
@@ 4.2 Implementation Details and Evaluation Metrics
In all experiments, the hyperparameters of DS-GIM (decay coefficient) and DMW-LA (linear function) remained unchanged. For different datasets, just set k in K-means as the number of categories in the dataset. The input resolutions of two encoders are set to 224$\times$224 and 256$\times$256 respectively. In addition, we train the model using an AdamW optimizer with a learning rate of 0.0001, and a weight decay of 1e-4. The model is trained on a single RTX 3090 GPU with a batch size 20 using Dice ($L_{dice}$) and Cross-entropy ($L_{ce}$) losses for 300 epochs. The specific proportion of these two losses is $Loss=0.7\cdot L_{dice}+0.3\cdot L_{ce}$. Finally, all experiments are performed in the Python 3.10.0 and Torch 2.3.0 environment. Only use random rotation and flipping as data augmentation. We choose the Dice score (DSC) and 95% Hausdorff distance (HD95) as evaluation metrics. The HD95 represents the ${95}^{th}$ percentile distance between the boundaries of the model prediction graph and real labels.
@@ 4.3 Comparison with SOTA Methods
Visual Comparison and Analysis. Figure 6 shows the visualizations of Sim-MPNet and other networks on Synapse and SegPC-2021. Firstly, the segmentation results of Sim-MPNet are very close to the ground truth in Synapse. Especially when dealing with organs like the Pancreas, which have complex contours and rich details, Sim-MPNet shows better segmentation performance compared to other networks. Secondly, Sim-MPNet still obtains the best segmentation results for images containing multiple cells in SegPC-2021. Although false positives and false negatives may also occur, the incidence of these errors is significantly reduced compared to other networks. In addition, the semantic clustering receptive fields of DMW-LA on Synapse, ACDC and SegPC-2021 see appendix B.
@@
Table 3: Ablation study of DMW-LA on four datasets. We use different encoders to verify their effectiveness separately. The input resolution of each encoder is the same as that of the comparative experiment, which is 224$\times$224 (Encoder 1) and 256$\times$256 (Encoder 2) respectively.
@@ 4.4 Ablation Study
Effectiveness of DS-GIM and W-LD Update Strategy. We conducted ablation experiments on DS-GIM and W-LD update strategy. It can be seen from Table 4 that compared with the random update strategy, our W-LD update strategy can minimize redundant information and retain useful features. In addition, the remarkable performance improvement in the experiment further confirms the effectiveness of DS-GIM in capturing the internal differences of features.
Dataset ACDC SegPC-2021 ISIC-2018 Synapse
w/o W-LD 91.53 82.63 90.22 83.75
w/o DS-GIM 91.67 82.58 90.47 83.64
w/o W-LD&DS-GIM 91.02 82.23 90.07 83.39
Sim-MPNet 92.18 83.12 90.85 84.34
Table 4: Ablation study of DS-GIM and W-LD update strategy on four datasets. In w/o W-LD, we use a random update strategy to replace the W-LD update strategy. The figures in the table are DSC.
@@ Acknowledgements
Supplementary Material
Appendix A Setting about $K$ in W-LD Update Strategy
To accurately locate the category features of the input image, we propose the W-LD update strategy to dynamically update the prototype memory bank. The following is an introduction to the $K$ parameter in the W-LD update strategy.
Figure 7: The variation curve of $K$ on the Synapse dataset. $M$ is the feature dimension of similarity memory prior.
In W-LD update strategy, $K$ represents the update range of prototype memory bank. To prevent model training collapse, we add a restriction to $K$ ($\frac{M}{4}\leq K\leq\frac{3M}{4}$) to ensure effective updates to the prototype memory bank. It has been proven that our measure is effective, and there is no training collapse in all comparative experiments and ablation experiments. Figure 7 show the curve of $K$ in the training process. It can be seen that with the increase of epoch, the model gradually converges, and $K$ is stable around $\frac{M}{2}$. In the initial stage of training, $K$ fluctuates violently, indicating that the model is learning useful patterns in data, and both large and small initial $K$ may have negative effects. Thus, we set the initial value of $K$ to $\frac{M}{2}$ in all experiments to cooperate with model training.
Appendix B Additional Visualization Results
In the process of medical image segmentation, we assign a semantic cluster (similarity memory prior) to the segmentation target in each image to extract key categorical features. Figure 8 shows the semantic cluster receptive field of DMW-LA on three datasets. We can observe that different clusters encode different semantic concepts. In the Synapse dataset, Cluster 2, Cluster 4 and Cluster 6 tend to focus on the region where the organs are located. In the ACDC dataset, Cluster 1 tends to focus on the left ventricle, while Cluster 2 focuses more on the right ventricle. In the SegPC-2021 dataset, Cluster 1 tends to focus on cytoplasm, while Cluster 3 focuses more on tissue fluid and other cells in the background. This distribution indicates that when processing images, each semantic cluster focuses on specific category features, thereby helping the model identify key structures in medical images.
Figure 8: Visualization of semantic clusters (similarity memory priors) using DMW-LA on Synapse, ACDC, and SegPC-2021. We calculate the distribution of similarity memory prior in the prototype memory bank, and highlight their receptive fields with white boxes.
Appendix C Evaluation Metrics
To better evaluate the performance of the network, we choose Dice score (DSC) and 95% Hausdorff distance (HD95) as evaluation metrics, as shown in Equation 8 and Equation 9.
${DSC=\frac{(2*TP)}{(2*TP+FP+FN)}}$ (8)
[…]
$\max_{b\in B}\{\min_{a\in A}\{d(b,a)\}\}\}$
In the above formula, $TP$ represents foreground information is predicted as foreground, $FP$ represents background information is predicted as foreground, and $FN$ represents foreground information is predicted as background. The HD95 represents the ${95}^{th}$ percentile of distances between the boundaries of A and B. DSC is used to evaluate the degree of overlap, while HD95 is used to evaluate distance. The combination of them can comprehensively evaluate the performance of the model.